### v2

@@ Abstract
keywords:
Diffusion models, online reinforcement learning, maximum entropy reinforcement learning, soft actor-critic
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm for training intelligent agents to make decisions in complex control tasks [1, 2, 3, 4, 5]. Traditionally, RL focuses on maximizing the expected cumulative reward, where the agent selects actions that yield the highest return in each state [6]. However, this approach often overlooks the inherent uncertainty and variability of real-world environments, which can lead to suboptimal or overly deterministic policies. To address these limitations, Maximum Entropy Reinforcement Learning (MaxEnt RL) incorporates entropy maximization into the standard RL objective, encouraging exploration and improving robustness during policy learning [7, 8, 9].
The Soft Actor-Critic (SAC) algorithm [10] is an effective method for achieving the MaxEnt RL objective, which alternates between policy evaluation and policy improvement to progressively refine the policy. With high-capacity neural network approximators and suitable optimization techniques, SAC can provably converge to the optimal MaxEnt policy within the chosen policy set. The choice of policy representation in SAC is crucial, as it influences the exploration behavior during training and determines the proximity of the candidate policies to the optimal MaxEnt policy. In complex multi-goal RL tasks, where multiple feasible behavioral modes exist, the commonly used Gaussian policy typically explores only a single mode, which can cause the agent to get trapped in a local optimum and fail to approach the optimal MaxEnt policy that captures all possible behavioral modes.
In this paper, we propose using diffusion models [11, 12, 13, 14], a powerful generative model, as the policy representation within the SAC framework. This allows for the exploration of all promising behavioral modes and facilitates convergence to the optimal MaxEnt policy. Diffusion models transform the original data distribution into a tractable Gaussian by progressively adding Gaussian noise, which is known as the forward diffusion process. After training a neural network to predict the noise added to the noisy samples, the original data can be recovered by solving the reverse diffusion process with the noise prediction network. While several generative models, e.g., variational autoencoders [15], generative adversarial networks [16], and normalizing flows [17] could serve as the policy representation, we choose diffusion models due to their balance between expressiveness and inference speed, achieving remarkable performance with affordable training and inference costs.
@@ 2.1 Maximum Entropy Reinforcement Learning
$J(\pi)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\sim\rho_{\pi}}\left[r(\boldsymbol{s}_{t},\boldsymbol{a}_{t})+\beta\mathcal{H}(\pi(\cdot|\boldsymbol{s}_{t}))\right],$ (1)
@@ 2.2 Soft Actor Critic
The optimal maximum entropy policy can be derived by applying the Soft Actor-Critic (SAC) algorithm [10]. In this subsection, we will briefly introduce the framework of SAC, and the relevant proofs are provided in Appendix A.1. The SAC algorithm utilizes two parameterized networks, $Q_{\theta}$ and $\pi_{\phi}$, to model the soft Q-function and the policy, where $\theta$ and $\phi$ represent the parameters of the respective networks. These networks are optimized by alternating between policy evaluation and policy improvement.
@@
Theorem 1 suggests that if the Bellman error can be reduced to zero and the policy loss is minimized at each optimization step, the soft actor-critic algorithm will converge to the optimal maximum entropy policy within the policy set $\Pi$. This indicates that the choice of the policy set $\Pi$ significantly affects the performance of the soft actor-critic algorithm. Specifically, a more expressive policy class will yield a policy closer to the optimal MaxEnt policy. Inspired by this intuition, we employ the diffusion model to represent the policy, as it is highly expressive and well-suited to capture the complex multimodal distribution [18, 19, 20, 21].
@@ 2.3 Diffusion Models
Diffusion models are powerful generative models. Given an unknown data distribution $p(\boldsymbol{x}_{0})$, which is typically a mixture of Dirac delta measures over the training dataset, diffusion models transform this data distribution into a tractable Gaussian distribution by progressively adding Gaussian noise [13]. In the context of a Variance-Preserving (VP) diffusion process [13, 14], the transition from the original sample $\boldsymbol{x}_{0}$ at time $t=0$ to the noisy sample $\boldsymbol{x}_{t}$ at time $t\in[0,1]$ follows the distribution:
@@
where $\alpha_{t}$ represents the log of the Signal-to-Noise Ratio (SNR) at time $t$, and $\sigma(\cdot)$ is the sigmoid function. $\alpha_{t}$ determines the amount of noise added at each time and is referred to as the noise schedule of a diffusion model. Denote the marginal distribution of $\boldsymbol{x}_{t}$ as $p(\boldsymbol{x}_{t})$. The noise schedule should be designed to ensure that $p(\boldsymbol{x}_{1}|\boldsymbol{x}_{0})\approx p(\boldsymbol{x}_{1})\approx\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{0},\boldsymbol{I})$, and that $\alpha_{t}$ is strictly decreasing w.r.t. $t$. Then, starting from $\boldsymbol{x}_{1}\sim\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{0},\boldsymbol{I})$, the original data samples can be recovered by reversing the diffusion process from $t=1$ to $t=0$. For sample generation, we can also employ the following probability flow ordinary differential equation (ODE) that shares the same marginal distribution with the diffusion process [14]:
@@
$L(\phi)$ $=\mathbb{E}_{t,\boldsymbol{x}_{t}}\left[w_{t}\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},\alpha_{t})+\sqrt{\sigma(-\alpha_{t})}\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t})\right\|^{2}_{2}\right]$ (6)
$=\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}\left[w_{t}\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},\alpha_{t})-\boldsymbol{\epsilon}\right\|^{2}_{2}\right]+C$ (7)
@@
Then we can solve the probability flow ODE in Equation 5 with the assistance of existing ODE solvers [13, 22, 23, 24, 25] to generate data samples.
@@ 3 Methodology
Consider a typical RL task that involves multiple behavior modes. The most efficient solution is to explore all behavior modes until one obviously outperforms the others. However, this exploration strategy is difficult to achieve with Gaussian policies. In the training process of a soft actor-critic algorithm with Gaussian policies, minimizing the K-L divergence between the Gaussian policy and the exponential of the Q-function—which is often multimodal in multi-goal tasks—tends to push the Gaussian policy to allocate most of the probability mass to the action region with the highest Q value [26]. Consequently, other promising action regions with slightly lower Q values will be neglected, which may cause the agent to become stuck at a local optimal policy.
@@ 3.1 Q-weighted Noise Estimation
which is an expectation with respect to the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$, a.k.a. the reverse transition distribution of the diffusion process. If samples from $p(\boldsymbol{a}_{0})$ are available, as is often the case in the application scenarios of diffusion models [27, 28, 18, 29, 30], we can first sample original actions $\boldsymbol{a}_{0}\sim p(\boldsymbol{a}_{0})$, and then sample noisy actions $\boldsymbol{a}_{t}\sim p(\boldsymbol{a}_{t}|\boldsymbol{a}_{0})$ to obtain several sample pairs following the joint distribution $p(\boldsymbol{a}_{0},\boldsymbol{a}_{t})$. Then for a fixed noisy action $\boldsymbol{a}_{t}$, the corresponding $\boldsymbol{a}_{0}$ will conform the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$, which can serve as Monte Carlo samples to estimate the expectation in Equation 11. Conversely, in the context of the soft actor-critic algorithm, we lack samples from the target distribution $p(\boldsymbol{a}_{0})$ but instead have access to a Q-function. Therefore, we must establish the relationship between the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$ and the Q-function.
Lemma 2.
@@
$p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})\propto\exp(\frac{1}{\beta}Q(\boldsymbol{a}_{0}))\mathcal{N}(\boldsymbol{a}_{0}|\frac{1}{\sqrt{\sigma(\alpha_{t})}}\boldsymbol{a}_{t},\frac{\sigma(-\alpha_{t})}{\sigma(\alpha_{t})}\boldsymbol{I})$ (12)
The proof is provided in Appendix A.2. Lemma 2 demonstrates that the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$ can be seen as a Gaussian distribution of $\boldsymbol{a}_{0}$ weighted by the exponential of the Q-function. Sampling from the Gaussian distribution is straightforward, we can apply importance sampling [31] to estimate the expectation in Equation 11.
Theorem 3.
@@
The derivation is detailed in Appendix A.3. Although this importance sampling estimate is unbiased, it exhibits high variance when the variance of the Q-function is large. To address this issue, we employ the weighted importance sampling approach [31] to reduce variance and stabilize the training process.
Theorem 4.
@@
$\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})$ $\approx\frac{1}{\sqrt{\sigma(-\alpha_{t})}}\cdot\sum_{i=1}^{K}\frac{w(\boldsymbol{a}_{0}^{i})}{\sum_{j=1}^{K}w(\boldsymbol{a}_{0}^{j})}\boldsymbol{\epsilon}^{i}$ (14)
$=\frac{1}{\sqrt{\sigma(-\alpha_{t})}}\sum_{i=1}^{K}\textnormal{softmax}(\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{1:K}))_{i}\boldsymbol{\epsilon}^{i},$ (15)
@@
We briefly compare our method with two previous approaches that approximate the exponential of a given function $Q(a)$. The QSM method [32] estimates the score function as $\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})\approx\nabla_{\boldsymbol{a}_{t}}\ \frac{1}{\beta}Q(\boldsymbol{a}_{t})$. This approximation requires $p(\boldsymbol{a}_{t})\propto\exp(\frac{1}{\beta}Q(\boldsymbol{a}_{t}))$, which is true only when the time $t$ is close to 00. Therefore, the score function estimation in QSM is imprecise for most values of $t$. Another method iDEM [33] proposes $\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})\approx\frac{1}{\sqrt{\sigma(\alpha_{t})}}\sum_{i=1}^{K}\text{softmax}(\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{1:K}))_{i}\nabla_{\boldsymbol{a}_{0}^{i}}\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{i})$, and the derivation is included in Appendix A.4 for completion. Although the expressions of iDEM and our method appear similar and both can approach the true score function as $K\to\infty$, our method does not require computing the gradient of the Q-function, which is more computationally efficient, especially when the Q-function is evaluated on a neural network. Furthermore, the experiments in Section 5.2 demonstrate that the variance of the score estimation in our method is significantly lower than the other two methods that rely on gradient computation, leading to a more stable training process.
@@ 3.2 Probability Approximation of Diffusion Policy
Previous study [34, 35] proved that the log-likelihood of $p(\boldsymbol{a}_{0})$ can be written exactly as an expression that depends only on the true noise prediction target, i.e.,
$\log p(\boldsymbol{a}_{0})=c-\frac{1}{2}\int_{-\infty}^{+\infty}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\parallel\boldsymbol{\epsilon}-\boldsymbol{\epsilon}^{*}(\boldsymbol{a}_{t},\alpha_{t})\parallel^{2}_{2}\right]\text{d}\alpha_{t}$ (19)
@@
$\log p_{\phi}(\boldsymbol{a}_{0})=c-\frac{1}{2}\int_{-\infty}^{+\infty}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\parallel\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\boldsymbol{a}_{t},\alpha_{t})\parallel^{2}_{2}\right]\text{d}\alpha_{t}$ (20)
This corollary can be inferred from Equation 19. However, this expression is intractable because both the integral in $c$ and the integral of the noise prediction error diverge, with only their difference converging [34]. We attempt to approximate the integral using numerical integration techniques. However, we observe that using the log SNR as the integration variable results in a high variance, as it spans from $\infty$- ∞ to +$\infty$+ ∞. Therefore, we instead utilize $\sigma(\alpha_{t})$ with a narrower integration domain of $[0,1]$.
Theorem 6.

### v3

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm for training intelligent agents to make decisions in complex control tasks (Silver et al., 2016; Mnih et al., 2015; Kaufmann et al., 2023; Kiran et al., 2021; Ibarz et al., 2021). Traditionally, RL focuses on maximizing the expected cumulative reward, where the agent selects actions that yield the highest return in each state (Sutton & Barto, 1999). However, this approach often overlooks the inherent uncertainty and variability of real-world environments, which can lead to suboptimal or overly deterministic policies. To address these limitations, Maximum Entropy Reinforcement Learning (MaxEnt RL) incorporates entropy maximization into the standard RL objective, encouraging exploration and improving robustness during policy learning (Toussaint, 2009; Ziebart, 2010; Haarnoja et al., 2017).
The Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) is an effective method for achieving the MaxEnt RL objective, which alternates between policy evaluation and policy improvement to progressively refine the policy. With high-capacity neural network approximators and suitable optimization techniques, SAC can provably converge to the optimal MaxEnt policy within the chosen policy set. The choice of policy representation in SAC is crucial, as it influences the exploration behavior during training and determines the proximity of the candidate policies to the optimal MaxEnt policy. In complex multi-goal RL tasks, where multiple feasible behavioral modes exist, the commonly used Gaussian policy typically explores only a single mode, which can cause the agent to get trapped in a local optimum and fail to approach the optimal MaxEnt policy that captures all possible behavioral modes.
In this paper, we propose using diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021b), a powerful generative model, as the policy representation within the SAC framework. This allows for the exploration of all promising behavioral modes and facilitates convergence to the optimal MaxEnt policy. Diffusion models transform the original data distribution into a tractable Gaussian by progressively adding Gaussian noise, which is known as the forward diffusion process. After training a neural network to predict the noise added to the noisy samples, the original data can be recovered by solving the reverse diffusion process with the noise prediction network. While several generative models, e.g., variational autoencoders (Kingma, 2013), generative adversarial networks (Goodfellow et al., 2020), and normalizing flows (Rezende & Mohamed, 2015) could serve as the policy representation, we choose diffusion models due to their balance between expressiveness and inference speed, achieving remarkable performance with affordable training and inference costs.
@@ 2.1 Maximum Entropy Reinforcement Learning
$J(\pi)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\sim\rho_{\pi}}\left[r(\boldsymbol{s}_{t},\boldsymbol{a}_{t})+\beta\mathcal{H}(\pi(\cdot|\boldsymbol{s}_{t}))\right],$
@@ 2.2 Soft Actor Critic
The optimal maximum entropy policy can be derived by applying the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018). In this subsection, we will briefly introduce the framework of SAC, and the relevant proofs are provided in Appendix A.1. The SAC algorithm utilizes two parameterized networks, $Q_{\theta}$ and $\pi_{\phi}$, to model the soft Q-function and the policy, where $\theta$ and $\phi$ represent the parameters of the respective networks. These networks are optimized by alternating between policy evaluation and policy improvement.
@@
Theorem 2.1 suggests that if the Bellman error can be reduced to zero and the policy loss is minimized at each optimization step, the soft actor-critic algorithm will converge to the optimal maximum entropy policy within the policy set $\Pi$. This indicates that the choice of the policy set $\Pi$ significantly affects the performance of the soft actor-critic algorithm. Specifically, a more expressive policy class will yield a policy closer to the optimal MaxEnt policy. Inspired by this intuition, we employ the diffusion model to represent the policy, as it is highly expressive and well-suited to capture the complex multimodal distribution (Chi et al., 2023; Wang et al., 2023; Chen et al., 2023; Ajay et al., 2023).
@@ 2.3 Diffusion Models
Diffusion models are powerful generative models. Given an unknown data distribution $p(\boldsymbol{x}_{0})$, which is typically a mixture of Dirac delta measures over the training dataset, diffusion models transform this data distribution into a tractable Gaussian distribution by progressively adding Gaussian noise (Ho et al., 2020). In the context of a Variance-Preserving (VP) diffusion process (Ho et al., 2020; Song et al., 2021b), the transition from the original sample $\boldsymbol{x}_{0}$ at time $t=0$ to the noisy sample $\boldsymbol{x}_{t}$ at time $t\in[0,1]$ follows the distribution:
@@
where $\alpha_{t}$ represents the log of the Signal-to-Noise Ratio (SNR) at time $t$, and $\sigma(\cdot)$ is the sigmoid function. $\alpha_{t}$ determines the amount of noise added at each time and is referred to as the noise schedule of a diffusion model. Denote the marginal distribution of $\boldsymbol{x}_{t}$ as $p(\boldsymbol{x}_{t})$. The noise schedule should be designed to ensure that $p(\boldsymbol{x}_{1}|\boldsymbol{x}_{0})\approx p(\boldsymbol{x}_{1})\approx\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{0},\boldsymbol{I})$, and that $\alpha_{t}$ is strictly decreasing w.r.t. $t$. Then, starting from $\boldsymbol{x}_{1}\sim\mathcal{N}(\boldsymbol{x}_{1}|\boldsymbol{0},\boldsymbol{I})$, the original data samples can be recovered by reversing the diffusion process from $t=1$ to $t=0$. For sample generation, we can also employ the following probability flow ordinary differential equation (ODE) that shares the same marginal distribution with the diffusion process (Song et al., 2021b):
@@
$\small{L(\phi)}$ $\small{=\mathbb{E}_{t,\boldsymbol{x}_{t}}\left[w_{t}\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},\alpha_{t})+\sqrt{\sigma(-\alpha_{t})}\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t})\right\|^{2}_{2}\right]}$ (6)
$\small{=\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}\left[w_{t}\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},\alpha_{t})-\boldsymbol{\epsilon}\right\|^{2}_{2}\right]+C}$ (7)
@@
Then we can solve the probability flow ODE in Equation 5 with the assistance of existing ODE solvers (Ho et al., 2020; Song et al., 2021a; Lu et al., 2022; Karras et al., 2022; Zheng et al., 2023) to generate data samples.
@@ 3 Methodology
Consider a typical RL task that involves multiple behavior modes. The most efficient solution is to explore all behavior modes until one obviously outperforms the others. However, this exploration strategy is difficult to achieve with Gaussian policies. In the training process of a soft actor-critic algorithm with Gaussian policies, minimizing the KL divergence between the Gaussian policy and the exponential of the Q-function—which is often multimodal in multi-goal tasks—tends to push the Gaussian policy to allocate most of the probability mass to the action region with the highest Q value (Chen et al., 2024a). Consequently, other promising action regions with slightly lower Q values will be neglected, which may cause the agent to become stuck at a local optimal policy.
@@ 3.1 Q-weighted Noise Estimation
which is an expectation with respect to the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$, a.k.a. the reverse transition distribution of the diffusion process. If samples from $p(\boldsymbol{a}_{0})$ are available, as is often the case in the application scenarios of diffusion models (Saharia et al., 2022; Ho et al., 2022; Chi et al., 2023; Xu et al., 2023; Huang et al., 2023), we can first sample original actions $\boldsymbol{a}_{0}\sim p(\boldsymbol{a}_{0})$, and then sample noisy actions $\boldsymbol{a}_{t}\sim p(\boldsymbol{a}_{t}|\boldsymbol{a}_{0})$ to obtain several sample pairs following the joint distribution $p(\boldsymbol{a}_{0},\boldsymbol{a}_{t})$. Then for a fixed noisy action $\boldsymbol{a}_{t}$, the corresponding $\boldsymbol{a}_{0}$ will conform the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$, which can serve as Monte Carlo samples to estimate the expectation in Equation 11. Conversely, in the context of the soft actor-critic algorithm, we lack samples from the target distribution $p(\boldsymbol{a}_{0})$ but instead have access to a Q-function. Therefore, we must establish the relationship between the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$ and the Q-function.
Lemma 3.1.
@@
$p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})\propto\exp(\frac{1}{\beta}Q(\boldsymbol{a}_{0}))\mathcal{N}(\boldsymbol{a}_{0}|\frac{1}{\sqrt{\sigma(\alpha_{t})}}\boldsymbol{a}_{t},\frac{\sigma(-\alpha_{t})}{\sigma(\alpha_{t})}\boldsymbol{I})$
The proof is provided in Appendix A.2. Lemma 3.1 demonstrates that the conditional distribution $p(\boldsymbol{a}_{0}|\boldsymbol{a}_{t})$ can be seen as a Gaussian distribution of $\boldsymbol{a}_{0}$ weighted by the exponential of the Q-function. Sampling from the Gaussian distribution is straightforward, we can apply importance sampling (Bishop, 2006) to estimate the expectation in Equation 11.
Theorem 3.2.
@@
The derivation is detailed in Appendix A.3. Although this importance sampling estimate is unbiased, it exhibits high variance when the variance of the Q-function is large. To address this issue, we employ the weighted importance sampling approach (Bishop, 2006) to reduce variance and stabilize the training process.
Theorem 3.3.
@@
$\small{\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})}$ $\small{\approx\frac{1}{\sqrt{\sigma(-\alpha_{t})}}\cdot\sum_{i=1}^{K}\frac{w(\boldsymbol{a}_{0}^{i})}{\sum_{j=1}^{K}w(\boldsymbol{a}_{0}^{j})}\boldsymbol{\epsilon}^{i}}$ (14)
$\small{=\frac{1}{\sqrt{\sigma(-\alpha_{t})}}\sum_{i=1}^{K}\textnormal{softmax}(\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{1:K}))_{i}\boldsymbol{\epsilon}^{i},}$ (15)
@@
We briefly compare our method with two previous approaches that approximate the exponential of a given function $Q(a)$. The QSM method (Psenka et al., 2024) estimates the score function as $\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})\approx\nabla_{\boldsymbol{a}_{t}}\ \frac{1}{\beta}Q(\boldsymbol{a}_{t})$. This approximation requires $p(\boldsymbol{a}_{t})\propto\exp(\frac{1}{\beta}Q(\boldsymbol{a}_{t}))$, which is true only when the time $t$ is close to 00. Therefore, the score function estimation in QSM is imprecise for most values of $t$. Another method iDEM (Akhound-Sadegh et al., 2024) proposes $\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})\approx\frac{1}{\sqrt{\sigma(\alpha_{t})}}\sum_{i=1}^{K}\text{softmax}(\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{1:K}))_{i}\nabla_{\boldsymbol{a}_{0}^{i}}\frac{1}{\beta}Q(\boldsymbol{a}_{0}^{i})$, and the derivation is included in Appendix A.4 for completion. Although the expressions of iDEM and our method appear similar and both can approach the true score function as $K\to\infty$, our method does not require computing the gradient of the Q-function, which is more computationally efficient, especially when the Q-function is evaluated on a neural network. Furthermore, the experiments in Section 5.2 demonstrate that the variance of the score estimation in our method is significantly lower than the other two methods that rely on gradient computation, leading to a more stable training process.
@@ 3.2 Probability Approximation of Diffusion Policy
Previous study (Kong et al., 2023; Wu et al., 2024) proved that the log-likelihood of $p(\boldsymbol{a}_{0})$ can be written exactly as an expression that depends only on the true noise prediction target, i.e.,
$\log p(\boldsymbol{a}_{0})=c-\frac{1}{2}\int_{-\infty}^{+\infty}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\parallel\boldsymbol{\epsilon}-\boldsymbol{\epsilon}^{*}(\boldsymbol{a}_{t},\alpha_{t})\parallel^{2}_{2}\right]\text{d}\alpha_{t}$
@@
$\log p_{\phi}(\boldsymbol{a}_{0})=c-\frac{1}{2}\int_{-\infty}^{+\infty}\mathbb{E}_{\boldsymbol{\epsilon}}\left[\parallel\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\boldsymbol{a}_{t},\alpha_{t})\parallel^{2}_{2}\right]\text{d}\alpha_{t}$
This corollary can be inferred from Equation 19. However, this expression is intractable because both the integral in $c$ and the integral of the noise prediction error diverge, with only their difference converging (Kong et al., 2023). We attempt to approximate the integral using numerical integration techniques. However, we observe that using the log SNR as the integration variable results in a high variance, as it spans from $\infty$- ∞ to +$\infty$+ ∞. Therefore, we instead utilize $\sigma(\alpha_{t})$ with a narrower integration domain of $(0,1)$.
Theorem 3.5.
@@
$\log p_{\phi}(\boldsymbol{a}_{0})\approx c^{\prime}+\frac{1}{2}\sum_{i=1}^{T}w_{t_{i}}\left(d\cdot\sigma(\alpha_{t_{i}})-\tilde{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{a}_{t_{i}},\alpha_{t_{i}})\right)$